INFRARED PREDISSOCIATION SPECTROSCOPY OF CATIONIC ACETYLENE CLUSTERS, (C_{2}H_{2})_{n}^{+}, n = 1 − 4

\underline{RACHAEL A. RELPH}, JOSEPH C. BOPP, JOSEPH R. ROSCIOLI, MARK A. JOHNSON, Sterling Chemistry Laboratory, Yale University, PO Box 208107, New Haven, CT 06520.

Infrared predissociation spectroscopic studies of systematically solvated acetylene clusters, (C_{2}H_{2})_{n}^{+}, n = 1 − 4, reveal the structural evolution of these cations with increasing cluster size. The argon predissociation spectrum of the acetylene dimer coupled with harmonic frequency calculations suggests the dominate species adopts a cyclobutadiene-like geometry. The argon solvated trimer, (C_{6}H_{6})^{+} · Ar, predominantly loses (C_{2}H_{2})^{+} · Ar, suggesting that the trimer is composed of a covalently bonded “dimer core” which exhibits a spectrum different than cyclobutadiene, and is solvated by a weakly bound acetylene. The same spectral features are retained in the tetramer suggesting that the dimer core survives the solvation of additional acetylene molecules. The minor loss channel of (C_{6}H_{6})^{+} · Ar, loss of a single Ar, shows multiple features believed to be from a number of covalently bonded (C_{6}H_{6})^{+} species.